# Variational data assimilation in wall turbulence: From outer observations to wall stress and pressure

### v2

@@ 1 Introduction
Near-wall turbulence and the associated surface stresses are challenging to directly probe in experiments (Hutchins et al., 2009; Lee et al., 2016). The near-wall region is, however, instrumental to understanding the dynamics of wall turbulence, such as the dissipation of energy, generation of drag, and extreme stress events. In order to estimate the inner-layer turbulence from outer measurements, various reduced-complexity models have been proposed (Marusic et al., 2010; Illingworth et al., 2018; Sasaki et al., 2019). A recent study utilizing the Navier-Stokes equations demonstrated that turbulence in the near-wall region can be reconstructed with machine accuracy, when fully resolved measurements in the outer flow are imposed at every time step. This process, which is termed synchronization because the inner layer synchronizes to the true flow that generated the outer observations, is only possible if the first observation point is within approximately thirty viscous units from the wall. However, synchronization is possible only when the thickness of the unknown layer is within thirty viscous units. Beyond this critical thickness synchronization fails, and the highest accuracy of estimating the wall layer under such conditions has never been examined. This problem is investigated herein using four-dimensional adjoint-variational (4DVar) data assimilation. We report the optimal reconstruction of all the missing scales of near-wall turbulence, including the wall shear stresses and pressure, and the dependence of estimation accuracy on the available observations.
@@
Data assimilation techniques, especially when they strongly enforce the governing Navier-Stokes equations, are powerful tools for exploring the fundamental limitation of turbulence reconstruction (Zaki & Wang, 2021). The most intuitive approach to combine observations into equations is continuous data assimilation, which augments the equations with a forcing term that drives the estimated state towards observations (Lalescu et al., 2013; Clark Di Leoni et al., 2020). This approach has been adopted to identify the critical thickness of an unknown wall-attached horizontal layer that can be synchronized, given outer fully resolved velocity data (Wang & Zaki, 2022). If the amount of observations is insufficient for a perfect reconstruction, ensemble-variational methods (EnVar, Liu et al., 2008) or adjoint-variational approaches (4DVar, Dimet & Talagrand, 1986) are better alternatives to explore the highest possible estimation accuracy. Both methods aim at minimizing a cost function defined as the difference between observations and their estimation, with the evolution of estimated state being strictly required to satisfy the Navier-Stokes equations. The minimization of the cost function is distinct in ensemble and adjoint variational methods: The former approximates the gradient of the cost function using an ensemble of forward simulations (Mons et al., 2017; Buchta & Zaki, 2021); The latter computes the gradient by solving the adjoint equations (Zaki, 2024). When applied to reconstructing high-dimensional unknown parameters, such as initial or boundary conditions, the adjoint approach has been demonstrated to be more accurate and efficient than other data-assimilation methods (Mons et al., 2016). Wang & Zaki (2021) applied 4DVar to reconstruct turbulent channel flow from velocity data that were systematically sub-sampled in three-dimensional space and time, throughout the flow domain. The data resolutions were comparable to experimental measurements, and they demonstrated that the accuracy of the estimated flow was robust against measurement noise. In addition, the adjoint field provides the exact sensitivity of observations to the flow state, and thereby sheds light on the causality in turbulence dynamics. Therefore, we adopt an adjoint-variational approach to construct a Navier-Stokes solution that best reproduces velocity observations in the outer layer while generating all the missing scales in the wall layer. Such a solution, once obtained, provides a quantitative assessment of the optimal estimation of the near-wall turbulence from outer measurements. Another interpretation is that the assimilation identifies the missing portion of the near-wall flow that influences the outer observed scales. We will also explore the influence of Reynolds number and data resolution on the estimation accuracy.
@@ 2.2 Adjoint-variational data assimilation
$\mathcal{L}=\mathcal{J}+\left(\boldsymbol{u}^{{\dagger}},\frac{\partial\boldsymbol{u}}{\partial t}+\boldsymbol{u}\cdot\nabla\boldsymbol{u}+\nabla p-\frac{1}{Re}\nabla^{2}\boldsymbol{u}\right)+\left[p^{{\dagger}},\nabla\cdot\boldsymbol{u}\right],$ (3)
@@
$\nabla\cdot\boldsymbol{u}^{{\dagger}}$ $=$ $\frac{\mathcal{D}\mathcal{J}}{\mathcal{D}p},$ (5a)
$\frac{\partial\boldsymbol{u}^{{\dagger}}}{\partial t^{{\dagger}}}-\boldsymbol{u}\cdot\nabla\boldsymbol{u}^{{\dagger}}+(\nabla\boldsymbol{u})\cdot\boldsymbol{u}^{{\dagger}}$ $=$ $\nabla p^{{\dagger}}+\frac{1}{Re}\nabla^{2}\boldsymbol{u}^{{\dagger}}-\frac{\mathcal{D}\mathcal{J}}{\mathcal{D}\boldsymbol{u}}.$ (5b)
@@ 3.1 Performance of the algorithm: multiple windows and spectral analysis
Figure 7: ($a$) Streamwise velocity fluctuations in the adjoint-variational estimated field, true state, and linear stochastic estimation, at $t=6T$. The visualized fields are normalized by the true r.m.s. fluctuations at the corresponding wall-normal location. A version of this figure featured in the recent review by Zaki (2024). ($b$) Horizontally-averaged estimation error of streamwise velocity, reconstructed using (dashed) LSE and (solid) adjoint approach, normalized by the maximum value of r.m.s. streamwise fluctuations.
@@
Figure 9: Wall-normal distribution of the one-dimensional Fourier spectrum of estimation error, normalized by the spectrum of the true field. ($a$) Streamwise velocity, ($b$) spanwise velocity and ($c$) pressure. (i,ii) Streamwise and spanwise spectra. The thickness of the unknown layer is $l_{y}^{+}=50$. Figures ($ai$) and ($aii$) featured in the recent review by Zaki (2024).
@@ 3.2 The influence of the unobserved layer thickness
Figure 17: Estimation accuracy of wall signals versus thickness of the unknown layer, when fully resolved outer observations are adopted. The accuracy is quantified using the correlation coefficient between the true wall signals and their estimations. Symbols (,+,$\circ$,+$\times$ , + , ×) correspond to the two components of wall shear stress and pressure ($\tau_{xy},\tau_{yz},p$). A version of this figure featured in the recent review by Zaki (2024).
@@ 4 Conclusions
Funding. This work was supported in part by the Office of Naval Research (grant N00014-20-1-2715). Computational resources were provided by the Advanced Research Computing at Hopkins (ARCH) core facility (rockfish.jhu.edu), which is supported by the National Science Foundation (NSF) grant number OAC 1920103.

### v3

@@ 1 Introduction
Near-wall turbulence and the associated surface stresses are challenging to directly probe in experiments (Hutchins et al., 2009; Lee et al., 2016). The near-wall region is, however, instrumental to understanding the dynamics of wall turbulence, such as the dissipation of energy, generation of drag, and extreme stress events. In order to estimate the inner-layer turbulence from outer measurements, various reduced-complexity models have been proposed (Marusic et al., 2010; Illingworth et al., 2018; Sasaki et al., 2019). A recent study utilizing the Navier-Stokes equations demonstrated that turbulence in the near-wall region can be reconstructed with machine accuracy, when fully resolved measurements in the outer flow are imposed at every time step. This process is termed synchronization because the inner layer synchronizes to the true flow that generated the outer observations, and is only possible if the first observation point is within approximately thirty viscous units from the wall. Beyond this critical thickness synchronization fails, and the highest accuracy of estimating the wall layer under such conditions has never been examined. This problem is investigated herein using four-dimensional adjoint-variational (4DVar) data assimilation. We report the optimal reconstruction of all the missing scales of near-wall turbulence, including the wall shear stresses and pressure, and the dependence of estimation accuracy on the available observations.
@@
Data assimilation techniques, especially when they strongly enforce the governing Navier-Stokes equations, are powerful tools for exploring the fundamental limitation of turbulence reconstruction (Zaki & Wang, 2021). The most intuitive approach to combine observations into equations is continuous data assimilation, which augments the equations with a forcing term that drives the estimated state towards observations (Lalescu et al., 2013; Clark Di Leoni et al., 2020). This approach has been adopted to identify the critical thickness of an unknown wall-attached horizontal layer that can be synchronized, given outer fully resolved velocity data (Wang & Zaki, 2022). If the amount of observations is insufficient for a perfect reconstruction, ensemble-variational methods (EnVar, Liu et al., 2008) or adjoint-variational approaches (4DVar, Dimet & Talagrand, 1986) are better alternatives to explore the highest possible estimation accuracy. Both methods aim at minimizing a cost function defined as the difference between observations and their estimation, with the evolution of estimated state being strictly required to satisfy the Navier-Stokes equations. The minimization of the cost function is distinct in ensemble and adjoint variational methods: The former approximates the gradient of the cost function using an ensemble of forward simulations (Mons et al., 2017; Buchta & Zaki, 2021); The latter computes the gradient by solving the adjoint equations (Zaki, 2025). When applied to reconstructing high-dimensional unknown parameters, such as initial or boundary conditions, the adjoint approach has been demonstrated to be more accurate and efficient than other data-assimilation methods (Mons et al., 2016). Wang & Zaki (2021) applied 4DVar to reconstruct turbulent channel flow from velocity data that were systematically sub-sampled in three-dimensional space and time, throughout the flow domain. The data resolutions were comparable to experimental measurements, and they demonstrated that the accuracy of the estimated flow was robust against measurement noise. In addition, the adjoint field provides the exact sensitivity of observations to the flow state, and thereby sheds light on the causality in turbulence dynamics. Therefore, we adopt an adjoint-variational approach to construct a Navier-Stokes solution that best reproduces velocity observations in the outer layer while generating all the missing scales in the wall layer. Such a solution, once obtained, provides a quantitative assessment of the optimal estimation of the near-wall turbulence from outer measurements. Another interpretation is that the assimilation identifies the missing portion of the near-wall flow that influences the outer observed scales. We will also explore the influence of Reynolds number and data resolution on the estimation accuracy.
@@ 2.2 Adjoint-variational data assimilation
$\mathcal{L}=\mathcal{J}-\left(\boldsymbol{u}^{{\dagger}},\frac{\partial\boldsymbol{u}}{\partial t}+\boldsymbol{u}\cdot\nabla\boldsymbol{u}+\nabla p-\frac{1}{Re}\nabla^{2}\boldsymbol{u}\right)-\left(p^{{\dagger}},\nabla\cdot\boldsymbol{u}\right),$ (3)
@@
$\nabla\cdot\boldsymbol{u}^{{\dagger}}$ $=$ $-\frac{\mathcal{D}\mathcal{J}}{\mathcal{D}p},$ (5a)
$\frac{\partial\boldsymbol{u}^{{\dagger}}}{\partial t^{{\dagger}}}-\boldsymbol{u}\cdot\nabla\boldsymbol{u}^{{\dagger}}+(\nabla\boldsymbol{u})\cdot\boldsymbol{u}^{{\dagger}}$ $=$ $\nabla p^{{\dagger}}+\frac{1}{Re}\nabla^{2}\boldsymbol{u}^{{\dagger}}+\frac{\mathcal{D}\mathcal{J}}{\mathcal{D}\boldsymbol{u}}.$ (5b)
@@ 3.1 Performance of the algorithm: multiple windows and spectral analysis
Figure 7: ($a$) Streamwise velocity fluctuations in the adjoint-variational estimated field, true state, and linear stochastic estimation, at $t=6T$. The visualized fields are normalized by the true r.m.s. fluctuations at the corresponding wall-normal location. A version of this figure featured in the recent review by Zaki (2025, CC BY 4.0). ($b$) Horizontally-averaged estimation error of streamwise velocity, reconstructed using (dashed) LSE and (solid) adjoint approach, normalized by the maximum value of r.m.s. streamwise fluctuations.
@@
Figure 9: Wall-normal distribution of the one-dimensional Fourier spectrum of estimation error, normalized by the spectrum of the true field. ($a$) Streamwise velocity, ($b$) spanwise velocity and ($c$) pressure. (i,ii) Streamwise and spanwise spectra. The thickness of the unknown layer is $l_{y}^{+}=50$. Figures ($ai$) and ($aii$) featured in the recent review by Zaki (2025, CC BY 4.0).
@@ 3.2 The influence of the unobserved layer thickness
Figure 17: Estimation accuracy of wall signals versus thickness of the unknown layer, when fully resolved outer observations are adopted. The accuracy is quantified using the correlation coefficient between the true wall signals and their estimations. Symbols (,+,$\circ$,+$\times$ , + , ×) correspond to the two components of wall shear stress and pressure ($\tau_{xy},\tau_{yz},p$). A version of this figure featured in the recent review by Zaki (2025, CC BY 4.0).
@@ 4 Conclusions
Funding. This work was supported by the Office of Naval Research (N00014-20-1-2715).